OpenStudy (anonymous):

find the slope and the y intercept of the line 8-1/5y=2x

OpenStudy (mathteacher1729):

You want to use algebra to put the given equation into the form y = (some number)*x + (some other number) Do you know how to begin doing this?

OpenStudy (anonymous):

no clue a little confusing

OpenStudy (anonymous):

lizzy - is this 1/5 times y or 1 divided by 5y?

OpenStudy (anonymous):

1 divided by 5y

OpenStudy (anonymous):

wait 1/5 times y.

OpenStudy (anonymous):

ok

OpenStudy (mathteacher1729):

\[8-\frac{1}{5}y=2x\]

OpenStudy (anonymous):

yes that is it

OpenStudy (anonymous):

ok so we first need to multiply every term by 5 to get y: 5*8 - 5* (1/5) y = 5*2x 40 - y = 10x 40 - 10x = y or y = -10x + 40 now in this form we can pick out the slope and the intercept on y-axis

OpenStudy (anonymous):

-10 is the slope and y-axis intercept is the vwlue of y when x = 0 y = -10*0 + 40 y = 40

OpenStudy (anonymous):

did you understand the algebra?